# Adapting Large Language Models for Character-based Augmentative and Alternative Communication

### v2

@@ Acknowledgments
We thank the AAC user who provided their text for use in our final evaluation. This work was funded by the National Institutes of Health / National Institute on Deafness and Other Communication Disorders (R01DC009834) and by the National Science Foundation (IIS-1750193). Any opinions presented in this work are those of the authors and do not reflect the opinions of our funding agencies.
@@ A.1 Written test set
We obtained compositions by crowdsourced workers collected by Vertanen and Kristensson (2014). With the exception of a table of examples and the text from the AAC condition in Experiment 1, these novel compositions were never publicly available. Workers were asked to imagine they were writing on a mobile device and to invent a “fictitious but plausible message”. We used the data from condition Compose in Experiment 1, both conditions in Experiment 2, and the Compose condition in Experiment 3. See Vertanen and Kristensson (2014) for further details.
We reviewed workers’ text using a semi-automated process. We expanded abbreviations, corrected spelling and grammar errors, and split compositions into sentences. If we could not determine a worker’s intended text, we dropped the composition. We lowercased text and stripped end-of-sentence punctuation, dashes, and commas. We dropped sentences containing other characters (e.g. numbers). Table 4 shows some examples from the test set. We subsequently removed these sentences as well as those appearing in Vertanen and Kristensson (2014) from the test set. We also removed any sentences containing potentially identifiable information.
@@
We took only the unique compositions from a given worker. We only used workers who rated their English ability as native. We assigned each unique worker an anonymous number. A total of 227 unique workers were in our dataset. We made a development and test set from the even and odd numbered participants respectively. Our development and test sets contained 1,348 lines (8,645 words) and 1,347 lines (9,044 words) respectively.
@@ A.2 Spoken test set
Similar to past work Venkatagiri (1999); Vertanen (2013) we asked people to each write eight different questions or statements in response to an everyday communication situation (see Section A.4 for our exact task wording). We did this on Amazon Mechanical Turk. Each worker received one of the following situations:
@@
Our study was reviewed by our institutional review board (IRB) and judged to be exempt. Workers were paid $0.80 USD to complete the study. First workers completed a consent form and then created eight unique sentences. While we asked workers to rate their English ability, we found this rating was often not accurate. We reviewed each worker’s sentences and eliminated 54 workers who had poor English or misunderstood the task (e.g. talking about strategies for ending a conversation rather than what you might actually say).
Occasionally, we found workers wrote sentences that were conversational in nature, but did not match the target situation. We suspect this was due to worker inattention. We left these sentence in our test set since we felt they were representative of conversational messages an AAC user might want to write. This left us with 246 unique workers who were each assigned an anonymous number. We semi-automatically corrected obvious mistakes in spelling and grammar, and split compositions comprised of multiple sentences. We screened for and removed any sentences that contained potentially identifiable information.
Table 5 shows examples of sentences written in each situation. In general, we found workers’ compositions were creative and relevant to the target situation. After removing the sentences in Table 5, we made a dev and test set from the even and odd numbered participants respectively. Our dev and test sets contained 1,469 lines (11,261 words) and 1,360 lines (10,821 words) respectively.
@@ A.4 Crowdsourced task
Figure 1 shows the interface we used to collect AAC-like communications from workers on Amazon Mechanical Turk workers. We changed the text at the top based on the situation as follows:
@@ B.1 Sentence classifier hyperparameters
B.2 Subword hyperparameter tuning
Fine-tuning examples were single sentences in mixed case with punctuation and started with the model’s default start token of </s>. For C4 and Subtitle, we tuned using 4 M sentences chosen at random from the C4 0.90 and Subtitle 0.75 datasets. For Daily and BOLT we used all the training data from the datasets.
We tuned the following hyperparameters, allowing each to vary within the range given:
Dropout — Dropout rate during fine-tuning, [0.0, 0.75].
Weight decay — Weight decay during training, [1e-4, 1.0].
Batch size — Training batch size, 1120, 2240, 3360, 4480, 5600, 6720, 7840, or 8960. The batch size was chosen to maximize the memory utilization of an A100 GPU with 40 GB of memory and assuming final model training was done using distributed data parallel (DDP) processes on four GPUs.
Epochs — Training epochs over a dataset. We fixed the training epochs to 1 for C4 and Subtitle. We allowed to be in [1, 20] for BOLT and Daily.
@@ B.2 Subword hyperparameter tuning
| type | step | steps | rate | decay | size | trials | (ppl) | ||||
@@
Table 6 shows the tuned hyperparameter values as well as the average dev set perplexity for each step of our adaptation of opt-350m. Each step of the fine-tuning curriculum lowered the perplexity on the dev sets. However, we found simply tuning in one-step on a combination of BOLT and Daily yielded the best results, reducing the perplexity of the original opt-350m model from 2.30 to 2.07 (10% relative).
We compared the utility of each of our datasets in isolation in Table 7. We found BOLT performed the best, followed by Daily, Subtitle 0.75, C4 0.90, and finally Forum. To show that selecting sentences using our classification model helped, we fine-tuned on an equivalent random amount of C4 and Subtitle text. As shown in Table 7 (bottom), random data performed worse than using the threshold selected data. In the case of C4, random data degraded predictions compared to the base model.
@@ B.3 Potential subword improvements
As shown in Table 8, increasing or decreasing the threshold by 0.05 on C4 only increased perplexity. Subsequent adaptation of the C4 0.90 fine-tuned model on Subtitle also performed best (or the same) with the threshold of 0.75 versus thresholds of 0.70 or 0.80.
With recurrent neural network language models (RNNLMs), interpolating an RNNLM with an n-gram model as been shown to perform better than either model in isolation Mikolov et al. (2014). We computed the per-character probabilities for our spoken and written dev sets using 1) the n-gram optimized for AAC-like text, 2) the original opt-350m LLM, and 3) our best fine-tuned opt-350m LLM. Using SRILM Stolcke (2002), we found the optimal mixture weights between the n-gram and each LLM (the unadapted opt-350m and the fine-tune opt-350m). We found the optimal mixture weights with respect to the spoken and written dev sets independently.
For the unadapted LLM, the optimal mixture had a perplexity of 2.28, slightly lower than the 2.30 obtained using only opt-350m. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.12 and 0.29 respectively. However, the fine-tuned LLM mixture had the same perplexity of 2.10 using only the fine-tuned LLM. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.04 and 0.03 respectively. Thus it seems the gains offered by the n-gram model’s in-domain training data could be obtained instead by fine-tuning a pretrained LLM on in-domain data.
We tried starting adaptation examples with a unique three token sequence generated from the text “ AAC|” rather than the model’s default start token of “</s>”. We hoped the model would create an association between this token sequence and AAC-like text during training, and that this would encourage the model to generate more AAC-like text at inference time. We compared the opt-350m model fine-tuned in a single step on the combined BOLT and Daily data. We tuned separate model hyperparameters for each start token sequence. Using a starting context of “ AAC|” resulted in an average perplexity on our dev sets of 2.21. Using the default starting context of “</s>” resulted in a slightly lower average perplexity of 2.18.
@@ B.4 Byte model adaptation
| step | steps | rate | decay | size | trials | (ppl) | ||||
| ByGPT5 base | — | — | — | — | — | — | — | — | — | 2.86 |
| C4 0.90 | 1 | .458 | 760 | .096 | 2.77e-04 | .382 | 592 | 1 | 88 | 2.23 |
| Subtitle 0.75 | 2 | .433 | 55 | .948 | 6.10e-08 | .112 | 2368 | 1 | 156 | 2.25 |
| BOLT+Daily | 3 | .567 | 142 | .109 | 6.07e-06 | .109 | 6720 | 2 | 105 | 2.19 |
| BOLT+Daily | 1 | .701 | 423 | .421 | 4.83e-05 | .514 | 5600 | 10 | 77 | 2.31 |
As an alternative to a subword LLM and our search algorithm, we tested using the ByGPT5 LLM Belouadi and Eger (2023) that uses a byte tokenization instead. We fine-tuned the ByGPT5 Belouadi and Eger (2023) model similarly to the subword opt-350m model (Section B.2). Since fine-tuning took around four times as long compared to opt-350m, we used only 1 M sentences when tuning on C4 and Subtitle. This allowed the tuner to explore a similar number of parameter settings as opt-350m. For the byte model, the batch size was chosen from 592, 1184, 1776, 2368, 2960, 3552, 4144, or 4736.
Table 9 shows the tuned hyperparameters and the resulting dev set perplexity at each step of the adaptation curriculum. Predictions improved markedly compared to the base model likely due to the multilingual nature of the model. We found that unlike the subword opt-350m model, the byte model benefited from first fine-tuning on C4 and then Subtitle.
B.5 Classifier model adaptation
We also experimented with adding a classification layer to our domain-adapted version of opt-350m Zhang et al. (2022). Table 10 (rightmost column) shows the perplexity on the dev sets at each step in the curriculum. Because the base classification model had already received the domain adaptation, the primary aim of this curriculum was to teach the model the classification task. We did not observe large perplexity changes between any two stages as we did with the byte model, but the average perplexity did decrease slightly from step-to-step. Similar to during domain adaptation of opt-350m, we found combining BOLT and Daily in the final step improved performance. We found it helped to first train the classification model on C4 and Subtitle compared to training on just a combination of BOLT and Daily.
@@ B.5 Classifier model adaptation
| step | steps | rate | decay | size | sentence | trials | (ppl) | |||
@@
We fine-tuned the opt-350m Zhang et al. (2022) classifier model in the same way as the subword model (Section B.2). Since the training examples chose a location within each sentence (Section 6.3), we tuned the number of examples created from each sentence in place of epochs. For the C4 and Subtitle sets, this was fixed at 1. For BOLT, Daily, and BOLT+Daily, this was chosen from between 1 and 50, inclusive. Table 10 shows the selected hyperparameters for each step in the curriculum.
@@ Appendix C Search algorithm pseudocode
Our work in this paper relies heavily on the search algorithm described in Section 3. To help convey the structure of the algorithm more clearly, we produced pseudocode in addition to the Python code released in supplementary materials. Algorithm 1 describes how the vocab hash is built at initialization time to make the search more efficient. Since the search only needs to consider tokens that match the truncated portion of the context, we create a map from each possible sequence of text to the tokens that begin with that text. Algorithm 2 describes the search algorithm itself, which is run for each inference.

### v3

@@ Acknowledgments
We thank the AAC user who provided their text for use in our final evaluation. We thank Soufia Bahmani for her work on the subword LLM word prediction algorithm used in Table˜4. This work was funded by the National Institutes of Health / National Institute on Deafness and Other Communication Disorders (R01DC009834) and by the National Science Foundation (IIS-1750193 and IIS-2402876). Any opinions presented in this work are those of the authors and do not reflect the opinions of our funding agencies.
@@ A.1 Written test set
We obtained compositions by crowdsourced workers collected by Vertanen and Kristensson (2014). With the exception of a table of examples and the text from the AAC condition in Experiment 1, these novel compositions were never publicly available. Workers were asked to imagine they were writing on a mobile device and that they should invent a “fictitious but plausible message”. We used the data from condition Compose in Experiment 1, both conditions in Experiment 2, and the Compose condition in Experiment 3. See Vertanen and Kristensson (2014) for further details.
We reviewed workers’ text using a semi-automated process. We expanded abbreviations, corrected spelling and grammar errors, and split compositions into sentences. If we could not determine a worker’s intended text, we dropped the composition. Table˜5 shows some examples from the test set. We subsequently removed these sentences as well as those appearing in tables in Vertanen and Kristensson (2014) from the test set. We also removed any sentences containing potentially identifiable information.
@@
We took only the unique compositions from a given worker. We only used workers who rated their English ability as native. We assigned each unique worker an anonymous number. A total of 227 unique workers were in our dataset. We made a dev and test set from the even and odd numbered participants respectively. For the purposes of the experiments conducted here, we lowercased text and stripped end-of-sentence punctuation, dashes, and commas. We dropped sentences containing other characters (e.g. numbers). Our dev and test sets contained 1,348 lines (8,645 words) and 1,347 lines (9,044 words) respectively.
@@ A.2 Spoken test set
Similar to past work Venkatagiri (1999); Vertanen (2013), we asked people to each write eight different questions or statements in response to an everyday communication situation (see Section˜A.4 for our exact task wording). We did this on Amazon Mechanical Turk. Each worker received one of the following situations:
@@
Our study was reviewed by our institutional review board (IRB) and judged to be exempt. Workers were paid $0.80 USD to complete the study. First, workers completed a consent form and then created eight unique sentences. While we asked workers to rate their English ability, we found this rating was often not accurate. We reviewed each worker’s sentences and eliminated 54 workers who had poor English or misunderstood the task (e.g. talking about strategies for ending a conversation rather than what you might actually say). This left us with 246 unique workers who were each assigned an anonymous number.
Occasionally, we found workers wrote sentences that were conversational in nature, but did not match the target situation. We suspect this was due to worker inattention. We left these sentence in our test set since we felt they were representative of conversational messages an AAC user might want to write. We semi-automatically corrected obvious mistakes in spelling and grammar, and split compositions comprised of multiple sentences. We removed any sentences containing potentially identifiable information.
Table˜6 shows examples of sentences written in each situation. In general, we found workers’ compositions were creative and relevant to the target situation. After removing the sentences in Table˜6, we made a dev and test set from the even and odd numbered participants respectively. For the purposes of the experiments conducted here, we lowercased text and stripped end-of-sentence punctuation, dashes, and commas. We dropped sentences containing other characters (e.g. numbers). Our dev and test sets contained 1,469 lines (11,261 words) and 1,360 lines (10,821 words) respectively.
| Opening: |
@@ A.4 Crowdsourced task
Figure˜1 shows the web interface we used to collect AAC-like communications from workers on Amazon Mechanical Turk. We changed the text at the top based on the situation as follows:
@@ B.1 Sentence classifier hyperparameters
| type | steps | steps | rate | decay | size | trials | (ppl) | ||||
@@
B.2 Subword hyperparameter tuning
Fine-tuning examples were single sentences in mixed case with punctuation and started with the model’s default start token of </s>. For C4 and Subtitle, we tuned hyperparameters using four million sentences chosen at random from the C4 0.90 and Subtitle 0.75 datasets to allow a more thorough search of the parameter space in the allotted time. Once we selected hyperparameters, we trained the model using the entire dataset. For Daily and BOLT we used all the training data from the datasets for both hyperparameter tuning and training.
We tuned the following hyperparameters, allowing each to vary within the given ranges:
Dropout — Dropout rate during fine-tuning, [0.0, 0.75].
Weight decay — Weight decay during training, [1e-4, 1.0].
Batch size — Training batch size, 1120, 2240, 3360, 4480, 5600, 6720, 7840, or 8960. The batch size was chosen to maximize the memory utilization of an A100 GPU with 40 GB of memory and assuming final model training was done using distributed data parallel (DDP) on four GPUs.
Epochs — Training epochs over a dataset. We fixed the training epochs to 1 for C4 and Subtitle. We allowed the training epochs to be in [1, 20] for BOLT and Daily.
Table˜7 shows the tuned hyperparameter values as well as the average dev set perplexity for each step of our adaptation of opt-350m. Each step of the fine-tuning curriculum lowered the perplexity on the dev sets. However, we found simply tuning in one-step on a combination of BOLT and Daily yielded the best results, reducing the perplexity of the original opt-350m model from 2.30 to 2.07 (a 10% relative improvement).
We compared the utility of each of our datasets in isolation in Table˜8. We found BOLT performed the best, followed by Daily, Subtitle 0.75, C4 0.90, and finally Forum. To show that selecting sentences using our classification model helped, we fine-tuned on an equivalent random amount of C4 and Subtitle text. As shown in Table˜8 (bottom), random data performed worse than using the threshold selected data. In the case of C4, random data degraded predictions compared to the base model.
@@ B.3 Potential subword improvements
As shown in Table˜9, increasing or decreasing the threshold by 0.05 on C4 increased perplexity. Subsequent adaptation of the C4 0.90 fine-tuned model on Subtitle performed best (or the same) with a threshold of 0.75 versus 0.70 or 0.80.
With recurrent neural network language models (RNNLMs), interpolating an RNNLM with an n-gram model can perform better than either model in isolation Mikolov et al. (2014). We computed the per-character probabilities for our spoken and written dev sets using 1) the n-gram optimized for AAC-like text, 2) the original opt-350m LLM, and 3) our best fine-tuned opt-350m LLM. Using SRILM Stolcke (2002), we found the optimal mixture weights between the n-gram and each LLM (the unadapted and fine-tuned opt-350m models). We found the optimal mixture weights with respect to the spoken and written dev sets independently.
For the unadapted LLM, the optimal mixture had a perplexity of 2.28, slightly lower than the 2.30 obtained using only opt-350m. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.12 and 0.29 respectively. However, the fine-tuned LLM and n-gram mixture model had the same perplexity of 2.10 as using only the fine-tuned LLM. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.04 and 0.03 respectively. Thus it seems the gains offered by the n-gram model’s in-domain training data could be obtained instead by fine-tuning a pretrained LLM on in-domain data.
We tried starting each adaptation example with a unique three-token sequence generated from the text “ AAC|” rather than the model’s default start token of “</s>”. We hoped this would allow the model to associate this token sequence with the AAC-like text seen during fine-tuning. This might encourage the model to generate more AAC-like text when seeing this sequence again at inference time. We compared the opt-350m model fine-tuned in a single step on the combined BOLT and Daily data. We tuned separate model hyperparameters for each start token sequence. Using a starting context of “ AAC|” resulted in an average perplexity on our dev sets of 2.21. Using the default starting context of “</s>” resulted in a slightly lower average perplexity of 2.18.
@@ B.4 Byte model adaptation
| steps | steps | rate | decay | size | trials | (ppl) | ||||
| ByGPT5 base | — | — | — | — | — | — | — | — | — | 2.85 |
| C4 0.90 | 1 | .458 | 760 | .096 | 2.77e-04 | .382 | 592 | 1 | 88 | 2.21 |
| Subtitle 0.75 | 2 | .433 | 55 | .948 | 6.10e-08 | .112 | 2368 | 1 | 156 | 2.22 |
| BOLT+Daily | 3 | .567 | 142 | .109 | 6.07e-06 | .109 | 6720 | 2 | 105 | 2.14 |
| BOLT+Daily | 1 | .701 | 423 | .421 | 4.83e-05 | .514 | 5600 | 10 | 77 | 2.27 |
@@
| steps | steps | rate | decay | size | sentence | trials | (ppl) | |||
@@
As an alternative to a subword LLM and our search algorithm, we tested using the ByGPT5 LLM Belouadi and Eger (2023). ByGPT5 uses byte instead of subword tokenization. We fine-tuned ByGPT5 similarly to the subword opt-350m model (see Section˜B.2). Since fine-tuning took around four times as long compared to opt-350m, we used only one million sentences when tuning on C4 and Subtitle. This allowed the tuner to explore a similar number of parameter settings as opt-350m. For the byte model, the batch size was chosen from 592, 1184, 1776, 2368, 2960, 3552, 4144, or 4736 to maximize memory use during model fine-tuning.
Table˜10 shows the tuned hyperparameters and the resulting dev set perplexity at each step of the adaptation curriculum. Predictions improved markedly compared to the base model, likely due to the multilingual nature of the original model. We found that unlike the subword opt-350m model, the byte model benefited from first fine-tuning on C4 and then subsequently fine-tuning on Subtitle.
B.5 Classifier model adaptation
We experimented with adding a classification layer to our domain-adapted version of opt-350m Zhang et al. (2022). Table˜11 (rightmost column) shows the perplexity on the dev sets at each step in the curriculum. Because the base classification model had already received the domain adaptation, the primary aim of this curriculum was to teach the model the classification task. We did not observe large perplexity changes between stages as with the byte model. However, perplexity did decrease slightly from step-to-step. Similar to during domain adaptation of opt-350m, we found combining BOLT and Daily in the final step improved performance. We found it helped to first train the classification model on C4 and then Subtitle compared to training on just a combination of BOLT and Daily.
We fine-tuned the opt-350m Zhang et al. (2022) classifier model in the same way as the subword model (Section˜B.2). Since the training examples chose a location within each sentence (Section˜6.3), we tuned the number of examples created from each sentence in place of epochs. For the C4 and Subtitle sets, this was fixed at 1. For BOLT, Daily, and BOLT+Daily, this was chosen from between 1 and 50, inclusive. Table˜11 shows the selected hyperparameters for each step in the curriculum.
@@ Appendix C Search algorithm pseudocode
This work relies heavily on our search algorithm (Section˜3). To convey the algorithm’s structure more clearly, we produced pseudocode in addition to the Python code in our supplementary materials. Algorithm 1 describes the vocab hash built at startup to make the search efficient. The hash maps each possible text sequence to valid matching tokens. Algorithm 2 describes the search algorithm used for each inference.